• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Meral Reyhan

    Log entry time 23:57:43 on August 11, 2006

    Entry number 179278

    keyword=Swing Shift/Run Summary

    Shift Summary
    Continuing Kinematics Point 17
    16:15- S0 in
    17:01- Lecroy HV error, MF3alarm, Right Arm Top
    17:10- 10 uA, requested
    18:00- So out, 20 uA, p=.4804 GeV/c, Kinematics Point 17A
    18:02- Half Wave Plate in
    21:12- Kinematics Point 18
    21:12- changed spectrometer the theta=15.26, p=0.7340 GeV/c,
    -doors=3"+3/4", cycled quads
    21:14- Optics target, 20uA
    21:30- rebooted coda in attempt to correct dead time problems, forgot to
    -reset after configuring
    21:45- rebooted coda properly
    21:55- 4cm dummy target
    22:11 15cm dummy target
    22:27 Ta, 30 uA (beam is actually around 27uA)
    22:58- LH2, radiator 5%, 12uA
    23:04- Rate check begins
    23:12- radiator out, 20uA
    23:30- no beam line on cryotarget heater chart for LD2
    23:42- reduced current to 8uA to get appropriate Trigger 3 rate
    23:50- finished Rate Check
    23:52- production runs
    Run Summary
    Run # # Events current
    2551/20909 1.9M 20 uA
    2552/20910 1.9M 20uA
    2553/20911 1.9M 20uA, S0 in
    2554/20912 1M 10uA

    2555/20913 1.9M 20uA, Kin 17A
    2556/20914 1.9M, 3.8M
    2557/20915 1.9M, 5.7M
    2558/20916 1.9M, 7.6M
    2559/20917 1.9M, 9.5M
    2560/20918 1.9M, 11.4M
    2561/20919 1.9M, 13.3M
    2562/20920 1.9M, 15.2M
    2563/20921 1.9M, 17.1M
    2564/20922 1.9M, 19M
    2565/20923 1M, 20M
    2566 junk
    2567/29024 .5M, 20uA, optics, Kin 18
    2668/29025 .4M, 20uA, 4cm dummy
    2669/29026 .4M, 20uA, 15cm dummy
    2670/29027 .1M, 30uA, Ta
    2671/29028 .3M, 12uA, LH2, 5%
    2672/29029 .1M, 20uA, LH2, out
    2673/29028 junk
    2674/29029 junk 20uA, LD2, out
    2675/29030 .5M, 20uA, LD2, out
    2676/29031 .66M 8uA, LD2, 5%
    2677/29032 8uA, LD2, 5%